# DIPLI: Deep Image Prior Lucky Imaging for Blind Astronomical Image Restoration

### v1

@@ IV-C Synthetic Data
TABLE II: DISTS and LPIPS metrics of the observation and ground truth compared to the corresponding values for synthetic data reconstructions obtained by DIPLI, RVRT, and DiffIR2VR-Zero, respectively.
IV-D Real Data
Figure 7: Real data reconstructions for different celestial objects. Insets are added to guide the eye.
After demonstrating the effectiveness of the DIPLI model on synthetic data, we conducted experiments with our real-world data, presenting the results in Figure 7. Due to the absence of ground truth images, we did not use PSNR and SSIM metrics. Instead, we relied on the expert evaluation. Unfortunately, metrics such as the energy of Laplacian or BRISQUE did not prove to be suitable for the quality evaluation of the reconstruction.
V Conclusion
In this paper, we proposed enhancements to the Deep Image Prior (DIP) model tailored for astrophotography applications, addressing the unique challenges posed by limited training data. We refined the DIP model to process multiple frames by incorporating the Back Projection method along with the TVNet model for optical flow computation. Additionally, we applied a simplified Markov approach utilizing Monte Carlo estimation and Langevin dynamics to ensure unbiased estimates with minimal variance.
These advancements collectively reduce noise learning and mitigate loss function fluctuations during training, enhancing the DIP model’s stability and performance. Experimental validation on real and synthetic image sets demonstrated that our modified DIP model outperforms, in most cases, all the evaluated models DIP, RVRT, and DiffIR2VR-Zero at least in one of the chosen metrics.
Our results highlight the robustness and effectiveness of the enhanced DIP model in producing high-quality images from low-quality observations, even in the challenging context of astrophotography.
@@ V Conclusion
In future work, it might be beneficial to explore more efficient architectures and techniques for image restoration in astrophotography. Although our modified DIP model shows substantial improvements, further optimization of the model for computational efficiency remains crucial. Furthermore, investigating other hybrid strategies and advanced deep learning techniques could further enhance the robustness and accuracy of astronomical image restoration. In addition, expanding the scope of training data through synthetic data generation or transfer learning could mitigate the limitations posed by the scarcity of astrophotography data. Lastly, exploring the application of our enhanced DIP model to other domains within scientific imaging could provide valuable insights and broaden the impact of our work.
@@ Acknowledgment
The corresponding author acknowledges the research support and funding provided by Alex Gerko. The authors also express their gratitude to Andrew McCarthy for providing the astronomical photo dataset.

### v2

@@ 4.4 Experiments and Results
Table 1: Comparison of reconstruction metrics for synthetic data. PSNR, SSIM, DISTS, and LPIPS are reported for the observation and ground truth, alongside results from DIPLI, RVRT, and DiffIR2VR-Zero. The naive approach corresponds to DIP with an early stopping point fixed for the entire dataset (470 iterations). Best and second-best results are highlighted in bold.
Real Data
In the absence of ground truth, full-reference metrics such as PSNR, SSIM, LPIPS, and DISTS, are inapplicable. Figure 7 therefore presents qualitative comparisons together with two no-reference indicators: Laplacian energy and BRISQUE [56]. Laplacian energy measures high-frequency content and is widely used in astronomical imaging as a sharpness or focus measure; it provides a reasonable sanity check but is noise-sensitive and cannot by itself distinguish genuine detail from hallucinated structure. BRISQUE, an NSS-based model trained on natural images, exhibits poor transferability to astronomical scenes and produced unstable rankings under distribution shift and compound distortions. Accordingly, the analysis prioritizes visual assessment (with limited expert inspection as a sanity check), treats Laplacian energy as auxiliary evidence, and de-prioritizes BRISQUE for this domain.
Qualitative assessment indicates that DIPLI produces clean reconstructions with strong detail retention, effective noise suppression, and no visually apparent method-induced artifacts. By contrast, RVRT+ frequently exhibits fine-grained speckle and residual ringing, while DiffIR2VR-Zero yields perceptually sharp results that nonetheless over-smooth regions and occasionally hallucinate texture near high-frequency structures. These observations align with the synthetic-data benchmarks with ground truth (Figure 6, Table 1).
Figure 7: Real-data reconstructions for different celestial objects. The LQ image pivot is shown on the left, and the remaining images present the corresponding HQ restorations. Insets are included to highlight specific details. Laplace Energy and BRISQUE metrics are reported. Higher values are better for Laplace Energy, and lower values are better for BRISQUE.
5 Conclusion
This work introduced DIPLI, an enhanced DIP-based framework designed for astrophotography applications. It addresses the unique challenges posed by limited training data by aggregating multiple frames via back-projection, aligning them using TVNet-based optical flow, and regularizing optimization through a Bayesian scheme based on stochastic gradient Langevin dynamics with Monte Carlo averaging.
Experiments on synthetic and real sequences indicate that DIPLI achieves superior performance relative to DIP, RVRT, and DiffIR2VR-Zero on most evaluated criteria. On synthetic benchmarks with ground truth, DIPLI achieves the strongest quantitative scores across the chosen metrics; on real scenes without references, qualitative analysis confirms improved detail preservation with reduced artifacts and noise.
@@ 5 Conclusion
In future work, it might be beneficial to explore more efficient architectures and techniques for image restoration in astrophotography. Although DIPLI demonstrates substantial improvements, further optimization of the model for computational efficiency remains crucial. Furthermore, investigating other hybrid strategies and advanced deep learning techniques could further enhance the robustness and accuracy of astronomical image restoration. In addition, expanding the scope of training data through synthetic data generation or transfer learning could mitigate the limitations posed by the scarcity of astrophotography data. Lastly, exploring the application of the proposed model to other domains within scientific imaging could provide valuable insights and broaden the impact of this work.
@@ Acknowledgment
The corresponding author acknowledges the research support provided by Alex Gerko. The authors also express their gratitude to Andrew McCarthy for providing the astronomical photo dataset.
Author contributions statement
Suraj Singh conceived the experiments, Suraj Singh and Anastasiia Batsheva conducted the experiments, Oleg Rogov and Ahmed Bouridane analysed the results. All authors reviewed the manuscript.